OpenStudy (anonymous):

The x-intercepts of a parabola are 5 and -7. What is the equation of the axis of symmetry?

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

how wud i get the ans

hero (hero):

(x-5)(x+7) = 0

hero (hero):

You can use that to at least get the parabola. After that, you have to complete the square

OpenStudy (anonymous):

how though

OpenStudy (anonymous):

can yu show me the steps please.

hero (hero):

x^2+2x-35 = 0 x^2+2x = 35 x^2+2x+1 = 35+1 (x+1)^2 = 36

hero (hero):

You really didn't need to do all of that however

OpenStudy (amistre64):

since the vertex is halfway between the intercepts: add them up and divide by 2

hero (hero):

amistre, I was getting to that

OpenStudy (anonymous):

so 7+5/2

OpenStudy (anonymous):

?

OpenStudy (amistre64):

there is no +5 in there :)

OpenStudy (amistre64):

but heros got it

OpenStudy (amistre64):

symetry means : same on both sides; so if your in the middle your symetric

hero (hero):

Actually, for that equation, I think they want you to write y = (x+1)^2 - 36

OpenStudy (amistre64):

nah, x=-1 is it

OpenStudy (amistre64):

the eq for the axis of sym

hero (hero):

Oops, well in that case (5-7)/2